determine how many solutions exist use either elimination or substitution to find the solutions if any y=2x+3 and y= -x-4
Answer by mananth(16946)

2 x + 1 y = 3 .............1
1 x + 1 y = -4 .............2
Eliminate y
multiply (1)by -1
Multiply (2) by 1
-2 x -1 y = -3
1 x + 1 y = -4
Add the two equations
-1 x = -7.00
/ -1
x = 7
plug value of x in (1)
2 x + 1 y = 3
14 + 1 y = 3
y = -11
(7,-11) is the solution


m.ananth@hotmail.ca